# Markus Dertwinkel-Kalt    Max R. P. Grossmann

### v1

@@ 2.2 Basic survey design
[…]
@@ 3.1 Data and summary statistics
Treatment: Info Full $0.013$ $-0.042$ $0.173$ $-1.004$
@@
Treatment: Info Strict $-0.002$ $-0.070$ $1.228$ $-1.143$
@@
Treatment: Info Lenient $0.010$ $-0.028$ $0.418$ $-1.432$
@@
Single-family house $0.024$ $-0.048$ $0.615$ $0.108$
@@ 3.3 Political priorities predict behaviors and attitudes toward the law
Treatment: Info Full $0.008$ $-0.001$ $0.722$ $-1.158$
@@
Treatment: Info Strict $-0.004$ $-0.058$ $1.382$ $-1.186$
@@
Treatment: Info Lenient $0.009$ $-0.025$ $0.447$ $-1.440$
@@
Treatment: Info Full $-0.032$ $-1.739^{*}$
@@
Treatment: Info Strict $-0.079$ $-1.416$
@@
Treatment: Info Lenient $-0.085$ $0.137$
@@
Belief is about proprietors $-0.111$
@@ Appendix D Additional tables and figures
Treatment: Info Lenient $-0.032$ $-0.033$
@@
Single-family house $0.025$ $-0.049$ $0.419$ $0.146$
@@
Single-family house $-0.052$ $0.115$ $4.099^{*}$ $0.225$
@@
Treatment: Info Full $0.759$ $0.261$ $-0.316$ $-1.187$ $-0.045$ $0.525$ $-0.048^{*}$
@@
Treatment: Info Strict $-0.195$ $-0.571$ $0.765$ $-0.670$ $0.246$ $0.415$ $-0.013$
@@
Treatment: Info Lenient $0.000$ $0.196$ $-0.226$ $-0.129$ $0.743$ $-0.651$ $-0.003$
@@
Single-family house $-0.023$ $-0.893$ $-0.280$ $-0.496$ $1.167^{***}$ $0.381$ $-0.005$
@@
Treatment: Info Full $0.009$ $-0.010$ $0.672$ $-1.021$
@@
Treatment: Info Strict $-0.004$ $-0.041$ $1.568$ $-1.319$
@@
Treatment: Info Lenient $0.011$ $-0.024$ $0.560$ $-1.587$

### v2

@@ 2.2 Basic survey design
Startedn=8,542n=8{,}542Proprietor or landlordn=3,460n=3{,}4601. Demographics2. Information provision3. Attitudes4. Quiz5. Beliefs, motives,dictator game6. Policy preferencesCompleted experimentn=2,995n=2{,}995Passed attention checkn=2,668n=2{,}668
@@ 3.1 Data and summary statistics
Treatment: Info Full $0.013$ 0.042-0.042 $0.173$ 1.004-1.004
@@
Treatment: Info Strict 0.002-0.002 0.070-0.070 $1.228$ 1.143-1.143
@@
Treatment: Info Lenient $0.010$ 0.028-0.028 $0.418$ 1.432-1.432
@@
Single-family house $0.024$ 0.048-0.048 $0.615$ $0.108$
@@ 3.3 Political priorities predict behaviors and attitudes toward the law
Treatment: Info Full $0.008$ 0.001-0.001 $0.722$ 1.158-1.158
@@
Treatment: Info Strict 0.004-0.004 0.058-0.058 $1.382$ 1.186-1.186
@@
Treatment: Info Lenient $0.009$ 0.025-0.025 $0.447$ 1.440-1.440
@@
Treatment: Info Full 0.032-0.032 $1.739^{*}$
@@
Treatment: Info Strict 0.079-0.079 1.416-1.416
@@
Treatment: Info Lenient 0.085-0.085 $0.137$
@@
Belief is about proprietors 0.111-0.111
@@ Appendix D Additional tables and figures
Treatment: Info Lenient 0.032-0.032 0.033-0.033
@@
Single-family house $0.025$ 0.049-0.049 $0.419$ $0.146$
@@
Single-family house 0.052-0.052 $0.115$ $4.099^{*}$ $0.225$
@@
Treatment: Info Full $0.759$ $0.261$ 0.316-0.316 1.187-1.187 0.045-0.045 $0.525$ $0.048^{*}$
@@
Treatment: Info Strict 0.195-0.195 0.571-0.571 $0.765$ 0.670-0.670 $0.246$ $0.415$ 0.013-0.013
@@
Treatment: Info Lenient $0.000$ $0.196$ 0.226-0.226 0.129-0.129 $0.743$ 0.651-0.651 0.003-0.003
@@
Single-family house 0.023-0.023 0.893-0.893 0.280-0.280 0.496-0.496 $1.167^{***}$ $0.381$ 0.005-0.005
@@
Treatment: Info Full $0.009$ 0.010-0.010 $0.672$ 1.021-1.021
@@
Treatment: Info Strict 0.004-0.004 0.041-0.041 $1.568$ 1.319-1.319
@@
Treatment: Info Lenient $0.011$ 0.024-0.024 $0.560$ 1.587-1.587